### v1

# Some Optimizers are More Equal: Understanding the Role of Optimizers in Group Fairness

Abstract
We study whether and how the choice of optimization algorithm can impact group fairness in deep neural networks. Through stochastic differential equation analysis of optimization dynamics in an analytically tractable setup, we demonstrate that the choice of optimization algorithm indeed influences fairness outcomes, particularly under severe imbalance. Furthermore, we show that when comparing two categories of optimizers, adaptive methods and stochastic methods, RMSProp (from the adaptive category) has a higher likelihood of converging to fairer minima than SGD (from the stochastic category). Building on this insight, we derive two new theoretical guarantees showing that, under appropriate conditions, RMSProp exhibits lower parameter bias and improved fairness in a single optimization step compared to SGD. We then validate these findings through extensive experiments on three publicly available datasets, namely CelebA, FairFace, and MS-COCO, across different tasks as facial expression recognition, gender classification, and multi-label classification, using various backbones. Considering multiple fairness definitions including equalized odds, equal opportunity, and demographic parity, adaptive optimizers like RMSProp and Adam consistently outperform SGD in terms of group fairness, while maintaining comparable predictive accuracy. Our results highlight the role of adaptive updates as a crucial yet overlooked mechanism for promoting fair outcomes. We release the source code at: https://anonymous.4open.science/??????.
1 Introduction
Machine learning is increasingly applied to a wide variety of domains, many of which have significant social implications [pmlr-v81-buolamwini18a, 10.1145/3494672]. These include, but are not limited to, decision-making systems, risk assessment models, recommendation engines, and personalized health interventions [HEGDE2020104492, rahimi2022quantum]. These systems influence people’s lives and can lead to unintended biases or disparities [10.1145/3457607], for instance by treating different groups of people differently. To address this, researchers have developed various techniques to ensure group fairness in machine learning algorithms [10.1145/3577190.3614141, hosseini2025faces].
Prior work has demonstrated that a variety of factors such as training data imbalance [10.1007/978-3-030-87199-4_39] and model hyper-parameters [NEURIPS2023_eb3c42dd] can impact the performance of deep neural networks in terms of group fairness and bias. In this paper, we pose the following question for the first time:
Does the choice of optimization algorithm impact group fairness in deep
neural networks? and how?
The limited theoretical and empirical analysis in this area leaves a major gap in our understanding of how optimizers impact the inherent performance of models in terms of bias. Consequently, we believe that this question can have significant implications in developing models that perform consistently across various groups, as optimization is a fundamental component of any deep learning model.
In this paper, we bridge this gap by first analyzing the impact of optimization in group fairness in a toy setup. This initial study confirms our hypothesis: that indeed the choice of optimizer impacts group fairness in downstream applications, particularly in the presence of severe bias in the training set. Next, we delve deep into the problem by proposing and proving two new theorems that demonstrate that among the two primary families of optimization algorithms, adaptive gradient methods such as Adam and RMSProp, and stochastic gradient methods such as SGD, the former is more likely to exhibit better group fairness. Next, we validate our theoretical findings by conducting extensive experiments on three publicly available datasets-MS-COCO [10.1007/978-3-319-10602-1_48], CelebA [liu2015faceattributes], and FairFace [karkkainenfairface] across three different tasks: multi-label classification, facial expression recognition, and gender classification. We carefully evaluate fairness using three widely adopted metrics: Equalized Odds, Equal Opportunity, and Demographic Parity. The results are strongly aligned with our theoretical analysis, consistently demonstrating that adaptive gradient methods achieve better fairness scores compared to stochastic gradient methods.
Our contributions in this paper are summarized as follows: (1) First, we provide a detailed mathematical analysis to study the impact of optimization algorithms on fairness in a toy setting. Here, we demonstrate that adaptive gradient algorithms like RMSProp have a higher probability of converging to fairer minima. (2) Next, we establish and prove two theorems showing that under appropriate conditions: (a) RMSProp provides fairer updates than SGD, and (b) the worst-case bias introduced by RMSProp in a single optimization step has an upper bound equal to that of SGD. These theoretical insights highlight key properties of optimization algorithms that contribute to their advantages in terms of fairness. (3) Through extensive experiments on multiple publicly available datasets and tasks, we empirically validate our theoretical findings. We systematically evaluate the fairness of models optimized using adaptive gradient methods and compare the outcomes to those obtained from stochastic gradient methods. Our results consistently show that adaptive gradient methods lead to fairer minima. (4) To enable rapid reproducibility and contribute to the area, we make our code publicly available at: https://anonymous.4open.science
2 Related work
Group fairness. Bias in machine learning models lead to discriminatory outcomes, and as these models are increasingly being deployed in real-life applications, such biases can imposes risk to the society [10.1145/3457607]. This risk is particularly important in high-stake applications, as hiring [10.1093/ojls/gqab006], healthcare [lett2023translating], etc. When models yield biased outcomes, they favor certain demographic groups, which reinforces the current social inequalities. As an example, facial expression recognition models show higher error rates for darker skin tones [10.1007/978-3-030-65414-6_35]. Such biases undermine trust in machine learning models and pose ethical as well as legal challenges. To analyze fairness in machine learning, researchers in this area often distinguish the notion of “group fairness” and “individual fairness” [10.1145/3457607]. Group fairness focuses on this notion that machine learning models treat different demographic groups, such as those defined by race, the same. On the other hand, individual fairness requires that similar individuals be treated similarly. We focus on group fairness in this paper, as it is more recognized in practical applications [10.1145/3630106.3659006].
One of the main sources of bias in machine learning models is data imbalance [deng2023fifa]. When certain demographic subgroups are underrepresented, models prioritize patterns from the majority group, leading to biased decision. To tackle this, a variety of methods have been proposed that can be categorized into 3 main groups [pmlr-v162-chai22a]: pre-process, in-process, and post-process. pre-process methods often apply some type of transformation to data to make the demographic information less recognizable to the model. As an example, [10.1145/3375627.3375865] proposes to augment the train set by synthetically generated samples, thereby ensuring an equal distribution for demographic subgroups. In-process methods directly modify the lerarning algorithm of the model. As an example, [10.1145/3577190.3614141] proposes a new loss function based of maximum mean discrepancy statistical distance to enhance fairness. Post-process methods intervenes after training is finished. For instance,  [NEURIPS2021_d9fea4ca] introduces a post-processing approach using graph Laplacian regularization to enforce individual fairness constraints while preserving accuracy. Nevertheless, in practice, these fairness enhancing methods are rarely used in real-world pipelines [Olulana_Cachel_Murai_Rundensteiner_2024]. One major limitation is their computational overhead or interference with the training phase. As an example, pre-processing methods require extensive data augmentation or transformations. Considering these challenges, an alternative approach to enhance fairness is to use components that are inherently present in every training. One of the fundamental building blocks of modern deep learning is optimization [NEURIPS2024_646d2edf]. Understanding the role of different optimization algorithms in fairness is crucial for practitioners seeking to develop fairer models without additional fairness-specific modifications.
Optimization algorithms. Optimization algorithms in deep learning generally fall into two main categories: SGD and adaptive gradient methods, such as Adam [kingma2014adam] and RMSProp [maunderstanding]. SGD updates model parameters using a constant learning rate [NEURIPS2024_ca98452d], while adaptive methods dynamically adjust per-parameter learning rates based on past gradients, often leading to faster convergence [ma2022qualitative].The impact of these 2 families of optimization algorithms on group fairness remains less understood. Nonetheless, several studies have explored their effects in other areas. For instance, [maunderstanding] empirically demonstrate that models trained with SGD are more robust to input perturbations than those trained with adaptive methods. Fairness and robustness are often competing objectives in deep learning [NEURIPS2022_a80ebbb4]. This hints that adaptive optimizers may yield better fairness compared to SGD. However, this relationship has been neither studied, nor understood.
3 Method
3.1 Preliminaries
Optimization algorithms. Stochastic gradient methods like SGD and adaptive gradient methods such as RMSProp and Adam are widely used for optimizing deep neural networks [NEURIPS2020_08fb104b]. Given a loss function $\mathcal{L}(w)$, SGD updates the model parameters $w$ at each iteration $k$ as:
$w_{k+1}=w_{k}-\eta\nabla\mathcal{L}(w_{k}),$ (1)
where $\eta$ denotes the learning rate. Adaptive gradient methods improve the speed of convergence by applying coordinate-wise scaling to the gradient [pan2022toward]. Specifically, RMSProp normalizes the gradient with an exponentially decaying average of its squared values:
$v_{k}=\gamma v_{k-1}+(1-\gamma)\nabla\mathcal{L}(w_{k})^{2},\quad w_{k+1}=w_{k}-\frac{\eta}{\sqrt{v_{k}+\epsilon}}\nabla\mathcal{L}(w_{k}),$ (2)
where $v_{k}$ is the moving average of squared gradients, $\gamma$ is the decay factor, and $\epsilon$ is a small constant added for numerical stability. On the other hand, Adam works by considering both the first and the second moments of gradients:
$m_{k}=\beta_{1}m_{k-1}+(1-\beta_{1})\nabla\mathcal{L}(w_{k}),$ (3)
$v_{k}=\beta_{2}v_{k-1}+(1-\beta_{2})\nabla\mathcal{L}(w_{k})^{2},$ (4)
$\hat{m}_{k}=\frac{m_{k}}{1-\beta_{1}^{k}},\quad\hat{v}_{k}=\frac{v_{k}}{1-\beta_{2}^{k}},$ (5)
$w_{k+1}=w_{k}-\frac{\eta}{\sqrt{\hat{v}_{k}}+\epsilon}\hat{m}_{k},$ (6)
where $\beta_{1}$ and $\beta_{2}$ are the first and second moment decay factors, respectively.
Stochastic Differential Equations (SDE). Given the interval $[0,T]$ where $T>0$, we have  [song2021scorebased]:
$dX_{t}=a(X(t),t)dt+b(X(t),t)dB(t),$ (7)
where $X_{t}\in R^{d}$ is a stochastic process (the solution of the SDE), $a:R^{d}\times[0,T]\rightarrow R^{d}$ is the drift term, $b:R^{d}\times[0,T]\rightarrow R^{d\times s}$ is the diffusion term, and $B(t)\in R^{s}$ is $s$-dimensional Brownian motion. Intuitively, SDEs extend ordinary differential equations by incorporating stochastic noise. Analyzing functions involving stochastic processes requires tools beyond classical calculus. A fundamental theorem in stochastic analysis is Itô’s lemma, extending the chain rule to the functions of stochastic processes [anresampling]. Please see the Appendix A for Itô’s lemma. Unlike ODEs, the solution of an SDE is a stochastic process $X_{t}$. Formally, the solution can be represented as:
[…]
The second integral, also known as the Itô’s integral, captures the stochastic contribution driven by Brownian motion. It is defined as:
[…]
where $\pi(n)$ denotes a sequence of $n$ partitions in $[0,t]$, and $\lim$ shows convergence in probability. In many cases, explicitly computing $X_{t}$ is intractable. Instead, it is often more practical to analyze the probability distribution of $X_{t}$. The Fokker-Planck equation, also known as the forward Kolmogorov equation, describes the time evolution of the probability density function $p(x,t)$ of $X_{t}$ [lai2023fp] as:
$\frac{\partial p(x,t)}{\partial t}=-\nabla\cdot(a(x,t)p(x,t))+\frac{1}{2}\nabla\cdot\left(\nabla\cdot(b(x,t)b(x,t)^{t}p(x,t))\right),$ (10)
[…]
SDE approximations for optimization algorithms. SDEs provide a powerful framework for analyzing the dynamics of discrete optimization algorithms such as SGD and RMSProp in a continuous setup. While direct analysis of discrete updates is often challenging [NEURIPS2021_69f62956], SDE approximations enable the study of optimization trajectories. In the context of optimization, the sources of stochasticity include mini-batch gradient noise which are variations due to data heterogeneity (e.g., differences in mini-batches drawn from different subpopulations), among others. To formalize this, we adopt the Noisy Gradient Oracle with Scale Parameter (NGOS) [NEURIPS2022_32ac7101], which models stochastic gradients as $g(w)=\nabla\mathcal{L}(w)+\Theta z$, where $\Theta$ is a noise scale parameter, and $z$ is a random variable that follows a distribution with covariance matrix $\Sigma(w)$ and mean of zero. This captures the inherent randomness in stochastic optimization. A well-behaved NGOS is defined as $\nabla\mathcal{L}(w)$ being Lipschitz and $\sqrt{\Sigma(w)}$ being bounded. See Appendix C for a detailed description of a well-behaved NGOS. Note that these assumptions are common and standard in prior works that study optimization using SDEs [li2017stochastic].
Under well-behaved NGOS, SGD can be approximated by the following SDE [li2017stochastic]:
$dW_{t}=-\nabla\mathcal{L}(W_{t})dt+(\eta\Sigma(W_{t}))^{1/2}dB(t),$ (11)
where $W_{t}$ represents an approximation of $w_{k}$ at discrete time steps $k\,\eta$. More recently, SDE approximation for RMSProp has been proposed as a coupled system of equations [NEURIPS2022_32ac7101]:
$dW_{t}=-P_{t}^{-1}\left(\nabla\mathcal{L}(W_{t})dt+\Theta\,\eta\,\Sigma^{1/2}(W_{t})dB(t)\right),\quad du_{t}=\frac{1-\gamma}{\eta^{2}}(diag(\Sigma(W_{t}))-u_{t})dt,$ (12)
where $P_{t}=\Theta\,\eta\,diag(u_{t})^{\frac{1}{2}}+\epsilon\,\eta\,I$ and $u_{t}$ is defined as $\frac{v_{k}}{\Theta^{2}}$. Note that these SDE approximations converge to $w_{k}$ in the weak sense, meaning that while individual trajectories of the discrete and continuous processes may differ, their probability distributions remain close. Please see Appendix D for a formal formulation of the convergence.
3.2 SDE analysis of optimizers
To illustrate the different behaviors of SGD and RMSProp in terms of group fairness, let’s consider a warm-up example. In this example, assume a training set consisting of two equally-sized subgroups, denoted as subgroups 0 and 1. That is, each group constitutes 50% of the total population. Please note this example can be generalized over more than 2 subgroups. For simplicity, assume each subgroup has its own loss function, denoted as:
$\mathcal{L}_{0}(w)=\frac{1}{2}(w-1)^{2},\quad\mathcal{L}_{1}(w)=\frac{1}{2}(w+1)^{2}.$ (13)
The objective is to minimize a population-level loss function defined as the weighted sum of the subgroup losses as:
$\mathcal{L}_{\text{pop}}(w)=0.5\mathcal{L}_{0}(w)+0.5\mathcal{L}_{1}(w),$ (14)
where the weights correspond to the proportion of each subgroup in the population. In this setup, the optimal solution for subgroup 0 is $w_{0}^{*}=1$, while the optimal solution for subgroup 1 is $w_{1}^{*}=-1$. However, The population’s optimal solution is obtained by minimizing $\mathcal{L}_{\text{pop}}(w)$, leading to $w_{\text{pop}}^{*}=0$. The following lemma shows that under the demographic parity definition of group fairness [10.1145/3457607], $w_{\text{pop}}^{*}=0$ is the fairest minima.
Lemma 1.
Let $\mathcal{L}_{0}(w)=\frac{1}{2}(w-1)^{2}$ and $\mathcal{L}_{1}(w)=\frac{1}{2}(w+1)^{2}$ be the loss functions for subgroups 0 and 1, respectively. Define the population loss as $\mathcal{L}_{\text{pop}}(w)=0.5\mathcal{L}_{0}(w)+0.5\mathcal{L}_{1}(w).$ Under the demographic parity definition of fairness, the fairest minimizer of $\mathcal{L}_{\text{pop}}(w)$ is $w_{\text{pop}}^{*}=0.$
Proof.
See Appendix E. ∎
In practice, we do not have access to the full population above, rather we have access to the set $\Omega$ of $N$ individual samples, sampled uniformly from the population. Additionally, set $\Omega$ may not contain an equal number of samples from subgroups 0 and 1, thus introducing bias. We assume that the probability of subgroup 0 in $\Omega$ is $p_{0}$ while the probability of subgroup 1 is $p_{1}=1-p_{0}$. For a given set $\Omega$, the empirical minimization problem used in practice is defined as: $\mathcal{L}_{emp}(w)=\frac{1}{N}\sum_{r\in\Omega}\mathcal{L}_{q_{r}}(w),$ where $q_{r}\in\{0,1\}$ corresponds to samples from subgroups 0 and 1. Note that as $N\rightarrow\infty$, $\mathcal{L}_{emp}\rightarrow\mathcal{L}_{pop}$. In min-batch training with batch size of 1, at each iteration, the optimizer processes a single individual from subgroup 0 with a probability of $p_{0}$ and from subgroup 1 with a probability $p_{1}$. In other words, with both SGD and RMSProp, the model receives $\nabla\mathcal{L}_{0}(w)$ with a probability of $p_{0}$ and $\nabla\mathcal{L}_{1}(w)$ with a probability of $p_{1}$, updating $w$ according to Eqs. 1 and 2, respectively. If $p_{0}>p_{1}$, the optimization trajectory is biased towards the group-specific minimum at $w=1$, and $w=-1$ otherwise. This demonstrates that disproportionate sampling of subgroups leads to unfair minima, which is particularly relevant in practical settings. The following theorem establishes that when the sampling bias $|p_{0}-p_{1}|$ exceeds a certain threshold, RMSProp has a higher probability of converging to $w^{*}_{pop}$ compared to SGD.
Theorem 1.
Let $p_{0},p_{1}\in(0,1)$ with $p_{0}+p_{1}=1$ be the subgroup sampling probabilities for the loss functions $\mathcal{L}_{0}(w)=\frac{1}{2}(w-1)^{2}$ and $\mathcal{L}_{1}(w)=\frac{1}{2}(w+1)^{2}$. Suppose we optimize the empirical objective $\mathcal{L}_{\emph{emp}}(w)\;=\;\frac{1}{N}\sum_{r\in\Omega}\mathcal{L}_{q_{r}}(w),$ where each sample $q_{r}\in\{0,1\}$ is drawn i.i.d. with probability $p_{0}$ for subgroup 0 or $p_{1}$ for subgroup 1. Consider mini-batch gradient updates of size $1$ (i.e., one sample per iteration) using SGD and RMSProp optimization algorithms. Then there exists a constant $\Delta(p_{1}p_{2},\eta)>0$ such that, whenever $|p_{0}-p_{1}|>\Delta(p_{1}p_{2},\eta)$, we have: $\frac{p_{rms}(w^{*}_{pop})}{p_{sgd}(w^{*}_{pop})}>1$, where $p_{rms}(w^{*}_{pop})$ and $p_{sgd}(w^{*}_{pop})$ are the probabilities of RMSProp and SGD converging to fair minima $w^{*}_{pop}=0$, respectively.
Proof.
As mentioned earlier, the gradient of $\mathcal{L}_{emp}$ is $\nabla\mathcal{L}_{0}(w)$ with a probability $p_{0}$ and $\nabla\mathcal{L}_{1}(w)$ with a probability of $p_{1}$. Thus, the covariance is $\Sigma(W_{t})=4p_{0}p_{1}$. This means $\Sigma(W_{t})$ is independent of $W_{t}$. When convergence occurs for RMSProp, $u_{t}$ defined in Eq. 12 converges to $\Sigma(W_{t})$. Thus, $u_{t}=4p_{0}p_{1}$. With this in mind, let’s derive the SDE for RMSProp:
$dW_{t}=\frac{-1}{\eta(2\Theta\sqrt{p_{0}p_{1}}+\epsilon)}(\nabla\mathcal{L}_{pop}(W_{t})dt+2\Theta\eta\sqrt{p_{0}p_{1}}dB_{t}).$ (15)
Here, $\epsilon$ is added for numerical stability, but it can be ignored as it is commonly done so in prior work [NEURIPS2022_32ac7101]. Next, let’s derive the SDE for SGD:
$dW_{t}=-\nabla\mathcal{L}_{pop}(W_{t})dt+2\sqrt{\eta}\sqrt{p_{0}p_{1}}dB(t).$ (16)
For analyzing the two SDEs above, we use the Fokker-Planck equations, presented in Eq. 10. When convergence occurs, $\frac{\partial p(w,t)}{\partial t}=0$. Thus, Fokker-Planck for RMSProp can be written as:
$\frac{\partial}{\partial w}(p(w)\times\frac{\nabla\mathcal{L}_{pop}(w)}{2\eta\Theta\sqrt{p_{0}p_{1}}})+\frac{1}{2}\frac{\partial^{2}p(w)}{\partial w^{2}}=0.$ (17)
Substituting $\nabla\mathcal{L}_{pop}=\frac{p_{0}}{2}(w-1)+\frac{p_{1}}{2}(w+1)$ in the equation above, we obtain:
$\frac{\partial}{\partial w}(p(w)\times\frac{p_{0}(w-1)+p_{1}(w+1)}{4\eta\Theta\sqrt{p_{0}p_{1}}})+\frac{1}{2}\frac{\partial^{2}p(w)}{\partial w^{2}}=0.$ (18)
Similarly, the Fokker-Planck equation for SGD can be written as:
$\frac{\partial}{\partial w}(p(w)\times\frac{p_{0}(w-1)+p_{1}(w+1)}{2})+2\eta p_{0}p_{1}\frac{\partial^{2}p(w)}{\partial w^{2}}=0.$ (19)
Eq. 18 is an analytically solvable ordinary differential equation, which can be written as:
$p_{rms}(w)=\sqrt{\frac{\kappa}{\pi}}exp\left(-\kappa(w-(p_{0}-p_{1}))^{2}\right),$ (20)
where $\kappa=\frac{1}{4\eta\Theta\sqrt{p_{0}p_{1}}}$. As a sanity check, we can see that the expectation of the distribution in Eq. 20 is equal to $p_{0}-p_{1}$. This means in the case of $p_{0}=p_{1}$, i.e. no sampling bias, the weight converges to $w^{*}=0$, the unbiased minimum. Similarly, we can write the analytical solution of Eq. 19 as:
$p_{sgd}(w)=\sqrt{\frac{\vartheta}{\pi}}exp\left(-\vartheta(w-(p_{0}-p_{1}))^{2}\right),$ (21)
where $\vartheta=\frac{1}{8\eta p_{0}p_{1}}$. As we use a batch size of 1, we can set $\Theta=1$ [NEURIPS2022_32ac7101]. Then, $1>=2\sqrt{p_{0}p_{1}}$ inequality holds. With this in mind, by substituting $w^{*}_{pop}=0$ in Eqs. 20 and 21, we can conclude $\frac{p_{rms}(w^{*}_{pop})}{p_{sgd}(w^{*}_{pop})}>1$ holds if and only if $|p_{0}-p_{1}|>\Delta(p_{1}p_{2},\eta)$, where $\Delta(p_{1}p_{2},\eta)=\sqrt{\frac{\frac{1}{2}\ln{\frac{\vartheta}{\kappa}}}{\vartheta-\kappa}}$. ∎
Figure 1: Percentage of 1000 runs converging within the fair neighborhood for SGD and RMSProp, under severe bias ($p_{0}=0.1$) and mild bias ($p_{0}=0.3$).
Now, let’s perform a quick simulation to verify our findings. We use the loss functions defined in Eq. 13 and set $p_{0}=0.1$ and $p_{1}=0.9$ to induce a severe sampling bias. We fix the learning rate to $\eta=0.1$ and apply SGD and RMSProp for 100 epochs. Each experiment is repeated 1000 times, and at every epoch we record the fraction of runs converging to the neighborhood of the fair optimum $w^{*}_{pop}$, defined by $|w-w^{*}_{pop}|<0.2$. The result is shown in Fig. 1 (left). As it is evident from this figure, approximately 10% of the runs under RMSProp lie in the neighborhood of $w^{*}_{pop}$, whereas SGD rarely converges to this fair region. Now, let’s induce a milder bias by setting $p_{0}=0.3$ and $p_{1}=0.7$ and run the same experiment. The results are illustrated in Fig. 1 (right). As shown in this figure, in the milder bias, the two optimizers behave more similarly, yet RMSProp attains a slightly higher fraction of solutions near $w^{*}_{pop}$. These observations align with our theoretical findings in Theorem 1. To show that our findings are true for a wider range of $\eta$ and alternative definitions of the fair neighborhood, we refer the reader to Appendix H for additional experiments.
3.3 Group fairness analysis
The detailed SDE analysis carried out in Subsection 3.2 highlights the distinct behaviors of SGD and RMSProp concerning group fairness. However, extending this analysis to higher dimensions, particularly through the Fokker-Planck Eq. 10, introduces significant complexity. To address this challenge, we present two theorems that leverage alternative mathematical tools beyond SDE, to study the impact of SGD and RMSProp on group fairness. First, we illustrate how adaptive optimizers like RMSProp suppress subgroup disparities at the level of parameter updates. The second theorem quantifies the consequence, showing that this mechanism translates into provably smaller fairness violations in terms of demographic parity.
Theorem 2.
Consider a population that consists of two subgroups with subgroup-specific loss functions $\mathcal{L}_{0}(w)$ and $\mathcal{L}_{1}(w)$, sampled with probabilities $p_{0}$ and $p_{1}$, respectively. Suppose a stochastic (online) training regime, in which each parameter update is computed from a sample drawn from one of the two subgroups. Suppose the gradients $\nabla\mathcal{L}_{0}(w_{k})$ and $\nabla\mathcal{L}_{1}(w_{k})$ are well-behaved isotropic NGOs, namely $\mathcal{N}(\mu_{0},\Theta_{0}^{2}I)$ and $\mathcal{N}(\mu_{1},\Theta_{1}^{2}I)$, respectively. Then, the difference in parameter updates between subgroups 0 and 1 under RMSProp has an upper bound given by the corresponding difference under SGD. Consequently, RMSProp offers fairer updates across subgroups.
Proof.
For SGD, the difference between the parameter updates across subgroups can be quantified as $||\nabla\mathcal{L}_{0}-\nabla\mathcal{L}_{1}||$. However, as the RMSProp scales the updates using exponentially decaying average, computing the difference is more complex. Let’s compute the expected value of $v_{k}$ defined in Eq. 2:
$\mathbb{E}[v_{k}]=\gamma^{k}v_{0}+(1-\gamma)\Sigma_{i=0}^{k-1}\gamma^{i}\mathbb{E}[\nabla\mathcal{L}_{i}^{2}].$ (22)
Given that $\mathcal{L}$ (loss function) is a discrete mixture of Gaussians, $\mathbb{E}[\nabla\mathcal{L}_{i}^{2}]$ can be calculated as:
$\mathbb{E}[\nabla\mathcal{L}_{i}^{2}]=p_{0}\left(\mu_{0}^{2}+\Theta_{0}^{2}\mathbf{1}\right)+p_{1}\left(\mu_{1}^{2}+\Theta_{1}^{2}\mathbf{1}\right),$ (23)
where $\mathbf{1}$ denotes a vector of all 1s. Substituting the equation above in Eq. 22 yields:
$\begin{split}\mathbb{E}[v_{k}]=\gamma^{k}v_{0}+(1-\gamma^{k})[p_{0}\left(\mu_{0}^{2}+\Theta_{0}^{2}\mathbf{1}\right)+p_{1}\left(\mu_{1}^{2}+\Theta_{1}^{2}\mathbf{1}\right)].\end{split}$ (24)
The equation above demonstrates that after enough steps, given that $\gamma<1$, $\mathbb{E}[v_{k}]$ converges to Eq. 23. Next, let’s analyze the variance of the $j$th coordinate of $v_{k}$, $v_{k_{j}}$, to see its spread. To this end, we compute the variance of $\nabla\mathcal{L}_{i_{j}}^{2}$ using the law of total variance as:
$var(\nabla\mathcal{L}_{i_{j}}^{2})=\mathbb{E}[var(\nabla\mathcal{L}_{i_{j}}^{2}|Z)]+var(\mathbb{E}[\nabla\mathcal{L}_{i_{j}}^{2}|Z]),$ (25)
where $Z$ is a random variable denoting the subgroup. Finally, assuming that (a) $\Theta_{0}^{2}$ and $\Theta_{1}^{2}$ are larger than $\mu_{0}^{2}$ and $\mu_{1}^{2}$, as shown experimentally in prior work [NEURIPS2022_32ac7101], and (b) there is bias in the training set (i.e., $p_{0}p_{1}$ is small), the variance $\operatorname{var}(v_{k_{j}})$ can be computed as $\mathcal{O}((1-\gamma)(\Theta_{0_{j}}^{4}+\Theta_{1_{j}}^{4}))$. For a large $\gamma$, typical in practical training, $var(v_{k_{j}})$ becomes negligible, and thus $v_{k}\rightarrow E[v_{k}]$. Using this, we can approximate the difference between the parameter updates across subgroups for RMSProp as $||D(\nabla\mathcal{L}_{0}-\nabla\mathcal{L}_{1})||$, where $D$ is a diagonal matrix, whose $jj$th element can be computed as: $D_{jj}=\frac{1}{\sqrt{p_{0}\left(\mu_{0_{j}}^{2}+\Theta_{0_{j}}^{2}\right)+p_{1}\left(\mu_{1_{j}}^{2}+\Theta_{1_{j}}^{2}\right)+\epsilon}}$. As $\Theta^{2}>\mu^{2}$ and in stochastic (online) training setup, we can set $\Theta^{2}\rightarrow 1$, then $D_{jj}<1$. In other words, the eigenvalues of matrix $D$ is less than 1, thus proving the theorem. ∎
Theorem 2 shows that RMSProp, unlike SGD, shrinks disparities between subgroups in the updates. In other words, RMSProp prevents large gradient disparities from dominating the training dynamics. Consequently, the final model is less likely to be pulled toward the subgroup with the larger gradients. Over many iterations, this can help prevent large performance gaps between subgroups. While this does not establish a universal fairness guarantee, it outlines why RMSProp’s dynamics yield fairer minima across subgroups. The following theorem analyzes fairness through the lens of demographic parity and establishes that, under appropriate conditions, in a single iteration, RMSProp’s worst-case increase in the gap of demographic parity has an upper-bound no greater than that of SGD.
Theorem 3.
Consider a population that consists of two subgroups with subgroup-specific loss functions $\mathcal{L}_{0}(w)$ and $\mathcal{L}_{1}(w)$, sampled with probabilities $p_{0}$ and $p_{1}$, respectively. Suppose a stochastic (online) training regime in which each parameter update is computed from a sample drawn from one of the two subgroups. Suppose that the gradients $\nabla\mathcal{L}_{0}(w_{k})$ and $\nabla\mathcal{L}_{1}(w_{k})$ are well-behaved isotropic NGOs, namely $\mathcal{N}(\mu_{0},\Theta_{0}^{2}I)$ and $\mathcal{N}(\mu_{1},\Theta_{1}^{2}I)$, respectively, with $\mu_{1}$ > $\mu_{0}$. Then, in expectation, the worst-case increase in the demographic parity gap after one iteration of RMSProp has an upper-bound no greater than the corresponding increase under SGD.
Proof.
As described in the proof of Lemma 1, the following can be used to approximate the gap in demographic parity: $\mathcal{F}(w)=|[\mathcal{L}(w)|z=0]-[\mathcal{L}(w)|z=1]|$. Let’s define the function $\Psi(w_{k})$ = $\mathcal{L}_{0}(w_{k})-\mathcal{L}_{1}(w_{k})$. Also, define $\varphi(w_{k})=\frac{L_{0}(w_{k})-L_{1}(w_{k})}{|L_{0}(w_{k})-L_{1}(w_{k})|}$. Let’s expand $\mathcal{F}(w_{k})$ around $w_{k}$ using Taylor series:
$\mathcal{F}(w_{k}+\delta)-\mathcal{F}(w_{k})=\nabla\mathcal{F}(w_{k})^{t}\delta.$ (26)
We can compute $\nabla\mathcal{F}(w_{k})$ as:
$\nabla\mathcal{F}(w_{k})=\varphi(w_{k})\cdot(\nabla\Psi(w_{k})).$ (27)
If in Eq. 26 we set $\delta-\eta\nabla\mathcal{L}(w_{k})$, given the updates of SGD in Eq. 1, we can re-write Eq. 26 as:
$\mathcal{F}(w_{{k+1}_{sgd}})-\mathcal{F}(w_{k})=-\eta\varphi(w_{k})\nabla\Psi(w_{k})^{t}\nabla\mathcal{L},$ (28)
where $\mathcal{F}(w_{{k+1}_{sgd}})$ is the value of function $\mathcal{F}()$ after one iteration using SGD. Using the proof of Theorem 2, we can perform a similar operation for RMSProp:
$\mathcal{F}(w_{{k+1}_{rms}})-\mathcal{F}(w_{k})=-\eta\varphi(w_{k})\nabla\Psi(w_{k})^{t}D\nabla\mathcal{L}.$ (29)
In the equation above, $\mathcal{F}(w_{{k+1}_{rms}})$ is the value of $\mathcal{F}()$ after one iteration using RMSProp, and matrix $D$ is a diagonal matrix defined in the proof of Theorem 2. As $\mu_{0}<\mu_{1}$, then the worst case increase in the gap of demographic parity for SGD occurs when $\nabla\Psi(w_{k})$ aligns with $\nabla\mathcal{L}$. Mathematically, the increase can be modeled as:
$|\mathcal{F}(w_{{k+1}_{sgd}})-\mathcal{F}(w_{k})|=\eta|\varphi(w_{k})\nabla\Psi(w_{k})^{t}\nabla\mathcal{L}|\leq\eta\|\nabla\Psi(w_{k})\|\cdot\|\nabla\mathcal{L}\|.$ (30)
Similarity, for RMSProp, we have:
$\begin{split}|\mathcal{F}(w_{{k+1}_{rms}})-\mathcal{F}(w_{k})|=\eta|\varphi(w_{k})\nabla\Psi(w_{k})^{t}D\nabla\mathcal{L}|\leq\eta\|D\nabla\Psi(w_{k})\|\cdot\|\nabla\mathcal{L}\|\\ <\eta\|\nabla\Psi(w_{k})\|\cdot\|\nabla\mathcal{L}\|.\end{split}$ (31)
In the equation above, as diagonal elements of $D$ are strictly less than 1, we have a strict inequality. Comparing Eqs. 30 and 31, we conclude the proof. ∎
This theorem establishes that the adaptive learning rate of RMSProp, which scales updates based on past squared gradients, helps mitigate demographic parity gaps that can occur during training, whereas SGD does not offer such a mechanism.
4 Experiments
Setup. We use three public datasets that are widely used in this area: CelebA [liu2015faceattributes], FairFace [karkkainenfairface], and MS-COCO [10.1007/978-3-319-10602-1_48]. CelebA and FairFace are datasets commonly used for group fairness analysis, while we also use a general-purpose dataset, MS-COCO, for further completeness. The details about these datasets are presented in Appendix I. The training protocol, backbones, and fairness measures used in our experiments are described in Appendix F and G, respectively.
Implementation details. We use PyTorch [paszke2019pytorch] and up to four Nvidia A100 GPUs. The hyperparameter values are provided in Appendix J, which have been optimized using Weights & Biases [wandb].
Results and analysis. We first calculate the fairness metrics for the ViT backbone across different datasets and different sensitive attributes (gender: G, race: R, and age: A). Fig. 2 presents the results for SGD, RMSProp, and Adam. Considering $F_{EOD}$, Adam consistently outperforms SGD across all scenarios, mostly by substantial margins. For instance, on CelebA dataset with gender as the sensitive attribute, Adam achieves an 8% improvement over SGD. Furthermore, RMSProp outperforms SGD in 4 out of 5 scenarios, and they both achieve zero $F_{EOD}$ for MS-COCO dataset. Given that MS-COCO contains some extremely underrepresented object categories, this result suggests that the disparity captured by $F_{EOD}$ is largely driven by rare categories receiving no correct classifications. The largest gap between RMSProp and SGD is observed on the FairFace dataset with race as the sensitive attribute, where RMSProp outperforms SGD by 9%. Another point to mention is that Adam and RMSProp yield comparable $F_{EOD}$ values, as expected. Considering $F_{EOP}$, both Adam and RMSProp outperform SGD in all 5 scenarios. To shed more light, the largest difference between the two is $6\%$ in the case of FairFace dataset with age as the sensitive attribute. Additionally, unlike SGD, both RMSProp and Adam achieve a near-perfect $F_{EOD}$ in the case of CelebA dataset with gender as the sensitive attribute, indicating their effectiveness in reaching fairer minima.
Figure 2: Fairness for ViT across different datasets, attributes (G: gender, A: age, R: race), and metrics.
Regarding $F_{DPA}$, Adam consistently outperforms SGD, while RMSProp surpasses SGD in 4 out of 5 scenarios. Both optimizers achieve zero $F_{DPA}$ on MS-COCO with gender as the sensitive attribute. Interestingly, the largest discrepancy between RMSProp and SGD occurs on the FairFace dataset with age as the sensitive attribute. This aligns with our theoretical insights from Theorem 1. Specifically, FairFace exhibits extreme imbalance, with a minority-to-all ratio of approximately 0.9%, whereas CelebA with gender as the sensitive attribute has a substantially higher ratio of 42%. These findings highlight that as dataset imbalance intensifies, the gap between adaptive optimizers and SGD becomes more evident. We refer the reader to Appendix K for the results on fairness metrics for other backbones across CelebA, FairFace, and MS-COCO datasets with different sensitive attributes. The results are well-aligned with the results above.
Table 1: Comparison of accuracy and F1 across different optimizers and datasets.
Dataset Accuracy F1 score
SGD RMSProp Adam SGD RMSProp Adam
CelebA 91.23 91.54 92.08 92.12 91.17 92.09
MS-COCO 89.62 89.71 90.03 68.35 71.03 74.10
FairFace 89.41 91.37 92.20 91.13 92.07 92.17
To demonstrate that our training is successful and the models are well-behaved, we evaluate the classification performance of ViT on the three datasets for facial expression recognition, multi-label classification, and gender classification, respectively. The results are presented in Table 1. The performance of the other models are presented in Appendix L. Here, we can confirm that in most cases, all three optimizers exhibit comparable accuracy and F1 scores across all tasks. For instance, on the CelebA dataset, SGD and RMSProp achieve an accuracy of 91%, while Adam attains 92%. This trend is consistent with previous findings in the prior work [maunderstanding]. As evident, the improved fairness observed with adaptive gradient algorithms cannot be solely attributed to their performance. For instance, on the CelebA dataset, the F1 score for SGD is higher than that of RMSProp, whereas on FairFace, RMSProp achieves a higher F1 score than SGD. However, in both cases, RMSProp demonstrates better fairness performance, as shown in Fig. 2. This further supports our theoretical conclusion that RMSProp is fairer than SGD, regardless of overall performance.
Table 2: p-values from the Wilcoxon test comparing SGD vs. RMSProp and SGD vs. Adam optimizers on the CelebA dataset for gender and age attributes.
Metric Gender Age
SGD-RMSProp SGD-Adam SGD-RMSProp SGD-Adam
$F_{EOD}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$F_{EOP}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-3}$
$F_{DPA}$ $2\times 10^{-3}$ $1\times 10^{-3}$ $7\times 10^{-3}$ $3\times 10^{-3}$
To rigorously assess the significance of fairness improvement of Adam and RMSProp over SGD, we conduct the following experiment. We train ViT on CelebA dataset 10 times using SGD, RMSProp, and Adam, while recording fairness metrics for both gender and age attributes in each run. To evaluate statistical significance, we perform a paired Wilcoxon signed-rank test [benavoli2014bayesian] on the fairness metrics and report the corresponding p-values in Table 2. As shown in the table, most scenarios yield p-values in the order of 0.001, indicating a statistically significant improvement in fairness for adaptive gradient optimizers compared to SGD. The worst p-value in the table corresponds to SGD vs. RMSProp on CelebA with age as the sensitive attribute. However, even this value remains statistically significant under the conventional threshold of 0.05 [NEURIPS2024_ce7984e3, paleja2021utility].
Figure 3: Difference of RMSProp and SGD’s fairness in different male-to-all ratios for CelebA dataset with gender as sensitive attribute.
Theorem 1 provides theoretical insight that under the demographic parity fairness criterion, RMSProp is more likely to converge to a fairer minimum than SGD in highly imbalanced datasets. Moreover, it predicts that as the dataset becomes more balanced, the fairness advantage of RMSProp over SGD gradually diminishes. To empirically validate this behavior, we conduct an experiment where we randomly downsample the ‘male’ subgroup in the CelebA dataset from its native male-to-all ratio of 42% down to 22% and 2%. We then train a model on each downsampled version using both RMSProp and SGD. We repeat this process five times while recording the values of $F_{DPA}$ and $F_{EOD}$. Fig. 3 shows the results, illustrating the absolute difference in $F_{DPA}$ and $F_{EOD}$ between SGD and RMSProp across different male-to-all ratios. As expected, the gap in demographic parity is maximum at the most severe imbalance level (2% male-to-all ratio). As the male-to-all ratio increases (indicating less imbalance), the absolute difference between the two optimizers gradually decreases, a trend that also holds for equalized odds. This empirical finding strongly aligns with the theoretical predictions in Theorem 1.
5 Conclusion
This paper provided both theoretical and empirical evidence that the choice of optimizer can significantly affect the group fairness of resulting deep learning models. Through establishing an analytically tractable setup and analyzing it via stochastic differential equations, we showed that SGD and RMSProp can converge to different minima with regards to fairness, with RMSProp more frequently reaching fairer optima. Next, we provided two theorems that showed 1) adaptive gradient methods (e.g., RMSProp) shrink subgroup-specific gradients more effectively, thereby reducing unfair updates compared to their SGD counterparts; 2) In a single optimization step, the worst case domographic disparity of RMSProp has the upper bound given by that of SGD. We validated these theoretical findings on three diverse datasets, namely CelebA, FairFace, and MS-COCO, and on multiple tasks. Across a range of group fairness criteria, i.e, equalized odds, equal opportunity, and demographic parity, adaptive optimizers consistently led to fairer solutions. We hope our work will encourage broader adoption of adaptive optimizers in fairness-critical domains, and we envision this research as a step toward developing more equitable deep learning models.
Acknowledgment
We would like to thank anonymous.
Appendix
Appendix A Itô’s Lemma
Itô’s lemma extends the chain rule of classical calculus to stochastic calculus, thereby accommodating functions that depend on Brownian motion $B(t)$. Suppose that the following stochastic differential equation (SDE) is given:
$dX(t)=M(t,X(t))dt+N(t,X(t))dB(t).$ (32)
In the equation above, X(t) is a stochastic process, M() and N() are the functions of $X(t)$ and $t$. Let’s define stochastic process $Y(t)$ by:
$Y(t)=Q(t,X(t)),$ (33)
where, $Q()$ is a function of stochastic process $X(t)$ and $t$. Itô’s lemma states that we can compute $dY(t)$ by using the following:
$\begin{split}dY(t)=\left(\frac{d}{dt}Q(t,X)+\frac{d}{dx}Q(t,X)\cdot M(t,X(t))+0.5\frac{d^{2}}{dx^{2}}Q(t,X)N^{2}(t,X(t))\right)dt+\\ \left(\frac{d}{dx}Q(t,X)\cdot N(t,X(t))\right)dB(t).\end{split}$ (34)
In the above equation, as the term $0.5\frac{d^{2}}{dx^{2}}Q(t,X)N^{2}(t,X(t))$ does not appear in classic calculus, it is often referred to as Itô’s correction term.
Appendix B Fokker-Planck equation
In this section, for completeness, we provide the derivation of fokker-planck equation from a given SDE. This is a well-known equation in the analysis of SDEs. Suppose the following SDE is given:
$dX(t)=M(t,X(t))dt+N(t,X(t))dB(t).$ (35)
Also, suppose that $Q(x)$ is a function which is 00 outside of a bounded interval. For simplicity, we use $X$ instead of $X(t)$. By Itô’s lemma, we have:
$dQ(X)=\left(M(t,X)Q^{\prime}(X)+0.5N^{2}(t,X)Q^{\prime\prime}(X)\right)dt+N(t,X)Q^{\prime}(X)dB(t).$ (36)
Let’s integrate the equation above from 0 to $t$ and then take the expected value from both sides:
[…]
Suppose $X$ is independent of $B$, then as $\mathbb{E}[dB(\tau)]=0$, the second integral above is zero. Next, let’s take the derivative with regards to $t$ from both sides:
$\frac{d}{dt}\mathbb{E}[Q(X)]=\mathbb{E}[\left(M(t,X)Q^{\prime}(X)+0.5N^{2}(t,X)Q^{\prime\prime}(X)\right)]$ (38)
To make the distribution of $X$ appear in the equations, let’s use the definition of $\mathbb{E}[.]$ in the equation above:
[…]
In the equation above, $p(t,x)$ is the distribution of stochastic process $X$. By applying integration by part to the integrals in the Eq. 39, and re-arranging the terms, we get:
[…]
As the equation above holds for all $Q(x)$, then the other factor must equal to zero. Thus, we have:
$\frac{\partial}{\partial t}p(t,x)+\frac{\partial}{\partial x}(p(t,x)M(t,x))-0.5\frac{\partial^{2}}{\partial x^{2}}(p(t,x)N^{2}(t,x))=0$ (41)
The equation above is known as Fokker-Planck equation.
Appendix C Well-behaved NGOS
Noisy Gradient Oracle with Scale Parameter (NGOS) models stochastic gradients as $g(w)=\nabla\mathcal{L}(w)+\Theta z$, where $\Theta$ is a noise scale parameter, and $z$ is a random variable that follows a distribution with covariance matrix $\Sigma(w)$ and a mean of zero. An NGOS is well-behaved if the conditions below are satisfied [NEURIPS2022_32ac7101]:
1. $\nabla\mathcal{L}(w)$ is Lipschitz.
2. Square root of covariance matrix is Lipschitz and bounded.
3. Both $\nabla\mathcal{L}(w)$ and $\Sigma(w)$ are differentiable and their partial derivatives up to 4-th order have polynomial growth.
4. Noise distribution in the NGOS must have low skewness. In other words, there must exist a function of polynomial growth such that $|\mathbb{E}[z^{3}]|$ is equal or less than that function divided by $\Theta$.
5. Noise distribution must have bounded moments. In other words, there must exist a constant $\kappa$ such that $\mathbb{E}[||z||^{2k}]^{\frac{1}{2k}}\leq\kappa(1+||w||)$.
Gaussian noise is a well-behaved NGOS. However, the noise in NGOS is not restricted to Gaussian. Particularly, if the noise distribution is not heavy-tailed, then it satisfies requirements 4 and 5. Whether or not the noise distribution in real-world applications is heavy-tailed is still an open problem. Nevertheless, several experimental results denote that the conditions described above are not too strong [NEURIPS2022_32ac7101].
Appendix D Approximation quality of SDE for SGD
The dynamics of SGD can be approximated by the following SDE [li2017stochastic]:
$dW_{t}=-\nabla\mathcal{L}(W_{t})dt+(\eta\Sigma(W_{t}))^{1/2}dB(t),$ (42)
where $W_{t}$ represents an approximation of $w_{k}$ at discrete time steps $k\,\eta$ and $\Sigma$ is the covariance matrix of the underlying NGOS. The noises that make the individual paths for the above SDE and the SGD are independent processes. Hence, SGD and its SDE approximation do not share the same sample paths; rather, they are close to each other in “distribution” or “weak sense” [li2017stochastic]. To formalize this mathematically, suppose $\Xi$ be the set of functions, $f(x)$, of polynomial growth, i.e. $\exists a,b\in\mathbb{R}^{+}$ such that $|f(x)|<a(1+|x|^{b})$. Then, SDE defined in Eq. 42 is the order $r$ weak approximation of SGD if for each function $f(x)$ in $\Xi$, there exists a constant $C$ such that for all $k=1\dots N$, $|\mathbb{E}[f(W_{k\eta})]-\mathbb{E}[f(w_{k})]|<C\eta^{r}$. This is a well-known result in analyzing optimization algorithms using SDEs, first time proposed in [li2017stochastic]. Similar theorems for RMSProp and Adam are available in [NEURIPS2022_32ac7101].
Appendix E Proof of Lemma 1
The demographic parity criterion for fairness between two subgroups is defined as:
$P(\hat{y}|z=0)=P(\hat{y}|z=1),$ (43)
where $z$ denotes the subgroup, and $\hat{y}$ represents the output of a model. The gap in demographic parity serves as a fairness measure [10.1145/3457607]. A common approach to quantifying this gap is through the zero-one loss difference [li2022achieving]: $|[\ell_{0/1}(w)|z=0]-[\ell_{0/1}(w)|z=1]|$. As established in prior work [li2022achieving], the zero-one loss can be approximated by the training loss, leading to the following fairness measure $\mathcal{F}(w)=|[\mathcal{L}(w)|z=0]-[\mathcal{L}(w)|z=1]|$. Substituting the given loss functions $\mathcal{L}_{0}(w)$ and $\mathcal{L}_{1}(w)$, we obtain $\mathcal{F}(w)=2|w|$. Since $\mathcal{F}(w)$ achieves its minimum at $w=0$, this proves that $w=0$ is the fairest minimizer.
Appendix F Training protocol
For facial expression recognition, we employ RetinaFace [deng2020retinaface] to detect and crop faces from images. The images in the FairFace dataset are already cropped by default. As part of the preprocessing pipeline for all tasks, we normalize the images and resize them to $224\times 224$. Data augmentation consists of random horizontal flipping and random rotation. Facial expression recognition and gender classification tasks are trained using the cross-entropy loss $\mathcal{L}_{ce}$, defined as:
$\mathcal{L}_{ce}=-\frac{1}{N}\sum_{j=1}^{N}\left[y_{j}\log(\hat{y}_{j})+(1-y_{j})\log(1-\hat{y}_{j})\right].$ (44)
For the multi-label classification task, given the class imbalance in MS-COCO, we employ the focal loss [lin2017focal] to mitigate overfitting to majority classes. It is formulated as:
$\mathcal{L}_{focal}=-\frac{1}{N}\sum_{j=1}^{N}[\xi y_{j}(1-\hat{y}_{j})^{\upsilon}\log(\hat{y}_{j})+(1-\xi)(1-y_{j})\hat{y}_{j}^{\upsilon}\log(1-\hat{y}_{j})],$ (45)
where $\xi$ is the balancing factor and $\upsilon$ is the focusing parameter. To optimize the learning rate during training, we leverage Bayesian optimization [NIPS2012_05311655] combined with the Hyperband algorithm [li2018hyperband]. The best-performing models are trained twice, and we report the average performance across these runs.
We experiment with multiple backbone architectures, including ResNet-18 and ResNet-50 [he2016deep], VGG-16 [simonyan2014very], and a recently proposed vision transformer (tiny ViT)[wu2022tinyvit]. These architectures have been widely used in studies analyzing the impact of optimization methods on different learning paradigms[maunderstanding], as well as in fairness-focused works [10.1007/978-3-030-65414-6_35]. ResNet models leverage residual connections to facilitate deep representation learning, while VGG-16 provides a structured CNN design with uniform filter sizes. The tiny ViT, a transformer-based model, processes images as patch embeddings, and showed state-of-the-art performance [wu2022tinyvit]. This diverse selection allows us to investigate fairness effects across both convolutional and transformer-based architectures.
Appendix G Fairness and performance measures
We report the performance across different tasks using accuracy and F1. We evaluate group fairness using three widely recognized fairness criteria [10.1145/3457607]: equalized odds, equal opportunity, and demographic parity. A predictor is fair under equalized odds definition if the true positive and false positive rates are equal across demographic subgroups. We measure the violation of this criterion as:
$F_{EOD}=\min_{i,j,q}[min(\frac{p(\hat{y}=c_{i}|y=c_{i},z_{j})}{p(\hat{y}=c_{i}|y=c_{i},z_{q})},\frac{p(\hat{y}=c_{i}|y\neq c_{i},z_{j})}{p(\hat{y}=c_{i}|y\neq c_{i},z_{q})})].$ (46)
In the equation above, $c_{i}$ demonstrates class $i$. A classifier satisfies equal opportunity if the true positive rate is equal across demographic subgroups. The related fairness measure is defined as:
$F_{EOP}=\min_{i,j}(\frac{\sum_{c=1}^{C}\text{{TPR}}_{i,c}}{\sum_{c=1}^{C}\text{{TPR}}_{j,c}}).$ (47)
Where $\text{{TPR}}_{i,c}$ denotes the true positive rate for i-th subgroup on class $c$. Finally, a predictor is said to be fair from demographic parity point of view if the predicted label distribution is independent of sensitive attributes. The corresponding measure is defined as:
$F_{DPA}=\min_{i,j,q}\frac{p(\hat{y}=c_{i}|z_{j})}{p(\hat{y}=c_{i}|z_{q})}.$ (48)
Regarding all three fairness metrics we use, higher values indicate better (fairer) outcomes.
Appendix H Additional results for Theorem 1
In this section, we extend our empirical analysis of Theorem 1 by evaluating our findings with broader choices of learning rates and alternative definitions of the fair neighborhood.
Theorem 1 establishes that for subgroup sampling probabilities $p_{0},p_{1}\in(0,1)$ with $p_{0}+p_{1}=1$, and loss functions $\mathcal{L}_{0}(w)=\frac{1}{2}(w-1)^{2}$ and $\mathcal{L}_{1}(w)=\frac{1}{2}(w+1)^{2}$, the optimization of the empirical loss function:
$\mathcal{L}_{emp}(w)\;=\;\frac{1}{N}\sum_{r\in\Omega}\mathcal{L}_{q_{r}}(w),$ (49)
where each sample $q_{r}\in{0,1}$ is drawn i.i.d. with probability $p_{0}$ for subgroup 0 or $p_{1}$ for subgroup 1, leads to an increased likelihood of RMSProp converging to the fair optimum $w^{*}_{pop}=0$ as the imbalance between $p_{0}$ and $p_{1}$ grows. To empirically validate these theoretical insights, we conducted simulations using a fixed learning rate of $\eta=0.1$, training for 1000 independent runs per experiment. The fraction of runs converging to the neighborhood of the fair optimum, defined as $|w-w^{*}_{pop}|<0.2$, was recorded at each epoch. We call $0.2$ the fairness threshold. To assess the robustness of our findings, we extend our analysis by examining a broader range of $\eta$ values and explore more restrictive and relaxed fairness thresholds. Table 3 summarizes the hyperparameter configurations used in our experiments. Fig.s 4 to 9 illustrate the convergence behavior of RMSProp and SGD under the specified configurations in the table.
The figures reveal a consistent trend in the convergence behavior of RMSProp and SGD. In the severely biased setting ($p_{0}=0.1$), RMSProp exhibits a higher likelihood of converging to the fair neighborhood, regardless of the learning rate. As the bias decreases ($p_{0}=0.3$), the behavior of RMSProp and SGD becomes more similar. These experimental findings are in line with theoretical findings in Theorem 1. Additionally, a more restrictive fairness threshold reduces the probability of convergence to the fair neighborhood, whereas a relaxed threshold increases it. This is expected, as stricter definitions impose tighter constraints on acceptable fair solutions, while relaxed criteria naturally allow a higher fraction of runs to qualify as fair.
Table 3: Hyperparameter configurations for the experimental scenarios.
Configuration $\eta_{RMSProp}$ $\eta_{SGD}$ Fairness Threshold
1 0.01 0.1 0.2
2 0.1 0.2 0.2
3 0.01 0.1 0.1
4 0.01 0.1 0.4
5 0.1 0.2 0.1
6 0.1 0.2 0.4
Figure 4: Convergence rates of RMSProp ($\eta=0.01$) and SGD ($\eta=0.1$) to the fair neighbourhood, defined by threshold of $0.2$.
Figure 5: Convergence rates of RMSProp ($\eta=0.1$) and SGD ($\eta=0.2$) to the fair neighbourhood, defined by threshold of $0.2$.
Figure 6: Convergence rates of RMSProp ($\eta=0.01$) and SGD ($\eta=0.1$) to the fair neighbourhood, defined by threshold of $0.1$.
Figure 7: Convergence rates of RMSProp ($\eta=0.01$) and SGD ($\eta=0.1$) to the fair neighbourhood, defined by threshold of $0.4$.
Figure 8: Convergence rates of RMSProp ($\eta=0.1$) and SGD ($\eta=0.2$) to the fair neighbourhood, defined by threshold of $0.1$.
Figure 9: Convergence rates of RMSProp ($\eta=0.1$) and SGD ($\eta=0.2$) to the fair neighbourhood, defined by threshold of $0.4$.
Appendix I Datasets
CelebA [liu2015faceattributes] is a large-scale and real-world dataset, which include 202,599 facial images from over 10,000 individuals. Each image is annotated with 40 distinct attributes, and the dataset is officially partitioned into training, validation, and test sets. In this study, we use gender (male vs. female) and age (young vs. old) as sensitive attributes, while performing facial expression recognition based on the “happy” attribute. Notably, the dataset exhibits an imbalance in demographic distribution, with the probability of an image depicting a male being 0.42 and the probability of depicting a young individual being 0.77. This skewed distribution indicates the presence of bias in the training set.
FairFace [karkkainenfairface] is a large-scale face attribute dataset, mainly designed to mitigate racial bias in facial analysis tasks. It includes 108,501 facial images sourced primarily from the YFCC-100M Flickr dataset [thomee2016yfcc100m]. Each image is annotated for gender, age (9 sub-groups), and race (7 sub-groups). In this study, we leverage FairFace for gender classification, treating age and race as sensitive attributes. Given its broad demographic representation, FairFace serves as a robust benchmark in fairness related tasks.
MS-COCO [10.1007/978-3-319-10602-1_48] is a large-scale, real-world benchmark dataset widely used for multi-label classification, object detection, and image captioning. It consists of over 330,000 images, including more than 200,000 labeled images spanning 80 object categories. In this study, we leverage MS-COCO for multi-label classification. Notably, the dataset does not provide explicit gender annotations. To infer gender labels, we follow a protocol established in previous works [zhao2017men], utilizing the five associated captions per image. Specifically, we retain images where at least one caption contains either the term “man” or “woman”. Additionally, we remove object categories that are not strongly associated with humans. Human-associated objects are defined as those appearing at least 100 times alongside either “man” or “woman” in the captions, resulting in a reduced set of 75 object categories. Fig. 10 provides sample captions and their corresponding inferred gender labels. It is worth mentioning that the probability of an image depicting a male is 0.71.
(a) 1. a woman prepares several sub sandwiches at a deli counter.
2. a person behind a display glass preparing food
3. a lady behind a sneeze guard making sub sandwiches.
4. a woman behind a deli counter making sub sandwiches.
5. there is a woman that is making sandwiches at a deli
(b) 1. a man in a yellow shirt holding a white dog and making a face
2. a man holding a hot dog bun beside his puppy.
3. a man putting a hot dog bun on a puppy and pretending to eat
4. a man holding a white puppy and a red leash.
5. a man who is pretending to bite a puppy.
Figure 10: Two example images from MS-COCO dataset along with their captions. The inferred gender labels from the captions for the left and right images are “woman” and “man”, respectively.
Appendix J Hyperparameters
We employ Bayesian Optimization [NIPS2012_05311655] with Hyperband [li2018hyperband] algorithm to systematically tune the learning rate across different optimizers. Hyperband efficiently allocates training resources to more promising setups and progressively eliminate underperforming ones. It evaluates training at specific checkpoints, called “brackets”. We set Hyperband brackets to [10, 20, 40], allowing enough time for each run to demonstrate its potential. Additionally, the learning rate is sampled using a log-uniform distribution within the following search spaces:
  • SGD: $[10^{-3},10^{-1}]$
  • RMSProp: $[10^{-5},10^{-2}]$
  • Adam: $[10^{-5},10^{-2}]$
All models are trained for maximum of 60 epochs with a batch size of 64 and weight decay of $10^{-4}$. We use Pytorch default values for optimizer specific hyperparameters:
  • RMSProp: $\gamma$ = 0.9
  • Adam: $\beta_{1}$ = 0.9, $\beta_{2}=0.999$
Additionally, we use “Reduce-on-Plateau” scheduler of Pytorch to decrease the learning rate during the training. if no improvement is observed for 5 epochs, then the learning rate is multiplied by 0.5.
Appendix K Additional fairness results across different backbones and datasets
In this section, we present fairness metrics for SGD, RMSProp, and Adam across ResNet-18, ResNet-50, and VGG-16 backbones, evaluated on multiple datasets with different sensitive attributes.
ResNet-18: Fig. 11 illustrates fairness metrics for the ResNet-18 backbone. Considering $F_{EOD}$, both Adam and RMSProp consistently outperform SGD across all datasets and sensitive attributes. The only exception is the MS-COCO dataset, where both RMSProp and SGD yield an $F_{EOD}$ of zero. This is due to MS-COCO’s 75 object classes, some of which are highly underrepresented, leading to an effective fairness score of zero. The largest difference in $F_{EOD}$ is observed on the CelebA dataset with age as the sensitive attribute, where RMSProp significantly surpasses SGD. Adam and RMSProp exhibit comparable performance, as expected.
For the $F_{EOP}$ metric, both RMSProp and Adam outperform SGD across all datasets. For example, in the FairFace dataset with age as the sensitive attribute, Adam achieves an $F_{EOP}$ of 80%, compared to 72% with SGD. Regarding $F_{DPA}$, Adam surpasses SGD in four out of five cases. The only exception is CelebA with age as the sensitive attribute, where both optimizers yield similar $F_{DPA}$. This aligns with Theorem 1, as CelebA is not sufficiently biased for adaptive gradient methods to demonstrate a significant fairness advantage.
ResNet-50: Fig. 12 presents results for the ResNet-50 backbone. For $F_{EOD}$, Adam consistently outperforms SGD, while RMSProp surpasses SGD in four out of five cases, with MS-COCO again being the exception. The difference between Adam and SGD reaches approximately 15% in CelebA when using age as the sensitive attribute. Similar trends emerge for $F_{EOP}$, where both Adam and RMSProp outperform SGD across all datasets. Finally, for $F_{DPA}$, more biased datasets, such as FairFace with age as the sensitive attribute, show a higher gap between adaptive gradient methods and SGD.
VGG-16: Fig. 13 shows fairness metrics for the VGG-16 backbone. As with the previous architectures, Adam outperforms SGD in all cases for $F_{EOD}$. For $F_{EOP}$, both Adam and RMSProp demonstrate superior fairness compared to SGD. In FairFace with age as the sensitive attribute, the difference between Adam and SGD reaches approximately 10%. For $F_{DPA}$, the most significant difference is again observed in FairFace with age as the sensitive attribute, which reinforces our theoretical findings.
Across all three fairness metrics, we observe consistent trends across multiple architectures, datasets, and problems, including gender classification, facial expression recognition, and multi-label classification. These extensive experiments provide strong empirical evidence that adaptive gradient methods, such as Adam and RMSProp, converge to fairer minima with higher probability compared to SGD.
Figure 11: Fairness metrics for ResNet-18 across different datasets and sensitive attributes.
Figure 12: Fairness metrics for ResNet-50 across different datasets and sensitive attributes.
Figure 13: Fairness metrics for VGG-16 across different datasets and sensitive attributes.
Appendix L Additional classification results across different backbones and datasets
Table 4 reports the accuracy and F1 scores of SGD, RMSProp, and Adam optimizers across ResNet-18, ResNet-50, and ViT backbones. We evaluate these optimizers on CelebA (facial expression recognition), MS-COCO (multi-label classification), and FairFace (gender classification) datasets. For CelebA, the performance of SGD, RMSProp, and Adam is comparable, with only minor variations. For instance, using the VGG-16 backbone, SGD achieves an accuracy of $92.22\%$, while RMSProp and Adam yield $92.37\%$ and $91.42\%$, respectively. Similarly, for MS-COCO and FairFace, the performance differences across the optimizers remain limited. The largest accuracy gap between SGD and adaptive optimizers occurs on MS-COCO with the VGG-16 backbone, reaching approximately $5.5\%$. However, the highest disparity in fairness metrics does not occur on MS-COCO but rather on FairFace dataset. These findings indicate that the differences between SGD and adaptive optimization algorithms cannot be solely attributed to their performance. Instead, as demonstrated in Theorems 13, the normalization behavior of these optimizers plays a crucial role in shaping their fairness properties.
Table 4: Performance of different backbones across datasets. Each cell reports three numbers corresponding to SGD, RMSProp, and Adam, respectively.
Dataset Backbones
ResNet18 ResNet50 VGG16
Acc. F1 Acc. F1 Acc. F1
CelebA 91.12, 92.21, 92.19 91.41, 92.76, 92.80 91.19, 92.60, 92.39 91.33, 92.09, 92.25 92.22, 92.37, 91.42 92.20, 92.28, 92.50
MS-COCO 88.31, 91.10, 90.42 66.96, 71.26, 69.37 89.15, 91.89, 91.70 67.68, 70.48, 70.16 89.48, 90.11, 91.57 66.47, 72.04, 73.20
FairFace 86.16, 90.36, 90.31 87.15, 88.71, 90.30 88.48, 90.58, 92.41 89.67, 90.77, 91.57 91.70, 91.81, 92.09 90.59, 91.66, 92.69
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly articulate the main research question, and the contributions are explicitly listed in a numbered format.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Each theorem provided in the paper clearly states its assumptions.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Each theorem is preceded by a clear statement of assumptions, followed by a complete and rigorous proof.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide all relevant information, including details on open-source datasets, hyperparameter values, and implementation. Additionally, the source code is made publicly available.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We use publicly available datasets and have released the source code with sufficient instructions to reproduce results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper includes all relevant experimental details, including hyperparameter values, optimizer types, and other training configurations.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report statistical significance using Wilcoxon signed-rank tests, as shown in Table 2.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The implementation details section includes information on the compute resources used.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have thoroughly reviewed the NeurIPS Code of Ethics and confirm that our research adheres to its principles.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The paper provides practical insights to practitioners into how optimizer choice affects group fairness. We discuss the implications in the paper.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not involve the release of high-risk models or data.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All datasets and external tools used in the paper are publicly available and properly cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We release the source code used in our experiments, accompanied by a README file that includes the guidelines.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve any crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve research with human subjects and therefore does not require IRB approval
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs were not used as part of the core methodology or any original component of the research.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.